Use the different of two squares formula to factor. Be sure to factor out any common factor
6. y ^2-49
8. 4x^2-25
Can yoyu help me factor these two questions using the two square formula to factor.

Answer by htmentor(1343)

Since 49 = 7^2, the factorization is (y+7)(y-7) (6.)
Since 4=2^2, 25=5^2, the factorization is (2x+5)(2x-5) (8.)